OpenStudy (anonymous):

Given that f(x) = 4x − 3 and g(x) = quantity of 2x minus 1, divided by 3 solve for g(f(2)).

OpenStudy (anonymous):

choices are: 1 3 9 15

OpenStudy (phi):

to find g(f(2)) first find f(2) f(x) is "short" for 4x-3 f(2) means use the formula 4x-3 but with x replaced with a 2 (and 4x means 4*x) can you find f(2) ?

OpenStudy (anonymous):

it would be 6... right?

OpenStudy (phi):

f(2) means 4*2 - 3 what do you get when you simplify that ?

OpenStudy (phi):

what is 2 times 4 ? and then subtract 3

OpenStudy (anonymous):

I meant 5 lol.

OpenStudy (phi):

ok, now you know f(2) is 5 that means g( f(2) ) is the same as g(5) now we do the same thing, but use the formula for g \[ g(x) = \frac{2x-1}{3} \] can you find g(5) ?

OpenStudy (phi):

remember g(5) means replace x with 5: \[ g(5) = \frac{2\cdot 5 - 1 }{3} \] what is 2 times 5 ? then subtract 1 that will be the "top" all divided by 3

OpenStudy (anonymous):

3

OpenStudy (phi):

yes, g(5) is (10-1)/3 = 9/3 = 3